OpenStudy (anonymous):

Medals given. Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). y = four thirdsx + 8 y = four thirdsx + 3 y = -three fourthsx + 8 y = -three fourthsx + 3

OpenStudy (abmon98):

4x - 3y = -12 Change into slope intercept form -3y=-4x-12 y=4/3x+4 parallel lines share the same gradient. y=mx+c m=4/3 (-3,4) y=mx+c 4/3*-3+c=4 -4+c=4 c=8 y=mx+c y=4/3x+8

OpenStudy (anonymous):

can I ask you one or two more?